OpenStudy (anonymous):

Use a graph of f or some other method to determine what, if any, value to assign to f(a) to make f continuous at x = a. f(x) = x/8x^2−x ; a=0

OpenStudy (nikvist):

\[f(x)=\frac{x}{8x^2-x}\quad\mbox{or}\quad f(x)=\frac{x}{8x^2}-x\]